Palencia Paints Corporation has a target capital structure of 25% debt and 75% common equity, with no preferred stock. Its before-tax cost of debt is 11%, and its marginal tax rate is 40%. The current stock price is P0 = $30.00. The last dividend was D0 = $3.00, and it is expected to grow at a 6% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations.
Answer #1

cost of common equity=(D1/Current price)+Growth rate

=(3*1.06)/30+0.06

=16.6%

After-tax cost of debt=11*(1-tax rate)

=11*(1-0.4)=6.6%

WACC=Respective costs*Respective weight

=(16.6*0.75)+(0.25*6.6)

=14.1%